# Worksworld: A Domain for Integrated Numeric Planning and Scheduling of Distributed Pipelined Workflows

### 12 Mar 2026 v1

@@ 3 Background
In our setting, the datasets are individual or groups of source data messages aggregated at our input data components. Benoit et al. 2013 provide a survey and taxonomy to classify pipelined workflow scheduling characteristics and specify complexity for scheduling linear task graphs. With their taxonomy, we classify our problem as scheduling pipelined workflows across related-heterogenous compute and storage resources at distributed sites connected by structured-heterogenous, bandwidth bounded, network links. We employ a multi-objective performance model that minimizes resource utilization (i.e. cost) while constraining latency (corresponding to makespan in single dataset DAG scheduling). However, our permanent scheduling of components to resources rather than repeatedly scheduling dataset execution tasks removes the temporal sequencing that drives the scheduling complexity in classical pipelined workflows. The hardness of Worksworld instead arises from multi-resource placement under network constraints, which proves structurally identical to the NP-hard virtual network embedding problem (Fischer et al. 2013).
@@ 5 Methodology
Table 1: Available predicates and numeric fluents in Worksworld.
@@ 9 Conclusion
We introduced Worksworld, a numeric domain for integrated planning and scheduling of permanent workflows. We showed empirically that a state-of-the-art planner can solve the joint planning and scheduling problem we defined in reasonable time and space for simple workflows on complex resource graphs. The domain and our other engineering artifacts are available publicly to benefit the AI planning community. Beyond future work mentioned earlier, we intend to explore leveraging state dependent action costs in the domain to permit greater cost penalty when scheduling on high utilization instances by normalizing work scheduled with our state-dependent resource_available fluent. These state dependent variables would also permit a more conservative resource consumption based on actual rates instead of max rates for data components. We believe increasing the types of DSIs modelled in Worksworld could permit interesting tradeoffs for a planner deciding to schedule workflow components with strict latency requirements on memory-based interfaces; latency tolerant workflows on disk-backed interfaces. We also believe our workflow and resource graph facilitate easy division of work and believe a hierarchy of planners could collaboratively solve the scheduling portion of the problem.

### 18 Mar 2026 v2

@@ 3 Background
In our setting, the datasets are individual or groups of source data messages aggregated at our input data components. Benoit et al. 2013 provide a survey and taxonomy to classify pipelined workflow scheduling characteristics and specify complexity for scheduling linear task graphs. With their taxonomy, we classify our problem as scheduling pipelined workflows across related-heterogenous compute and storage resources at distributed sites connected by structured-heterogenous, bandwidth bounded, network links. We employ a multi-objective performance model that minimizes resource utilization (i.e. cost) while constraining latency (corresponding to makespan in single dataset DAG scheduling). However, our permanent scheduling of components to resources rather than repeatedly scheduling dataset execution tasks removes the temporal sequencing that drives the scheduling complexity in classical pipelined workflows. The hardness of Worksworld instead arises from multi-resource placement under network constraints, which proves structurally similar to the NP-hard virtual network embedding problem (Fischer et al. 2013).
@@ 5 Methodology
Table 1: Available predicates and numeric fluents in Worksworld.
@@ 9 Conclusion
We introduced Worksworld, a numeric domain for integrated planning and scheduling of permanent workflows. We showed empirically that a state-of-the-art planner can solve the joint planning and scheduling problem we defined in reasonable time and space for simple workflows on complex resource graphs. The domain and our other engineering artifacts are available publicly to benefit the AI planning community. Beyond future work mentioned earlier, we intend to explore leveraging state-dependent action costs in the domain to permit greater cost penalty when scheduling on high-utilization instances. This will require normalizing work scheduled with our state-dependent resource_available fluent. Supporting state dependent action costs will also permit a more conservative cost estimate based on actual rates instead of max rates for workflow components. Second, we posit that increasing the types of DSIs modelled in Worksworld could permit interesting tradeoffs for a planner. For example, the planner may schedule latency-sensitive components on memory-based interfaces at higher cost, and latency-tolerant components on disk-backed interfaces at lower cost. Finally, our workflow and resource graph facilitate easy division of work that may enable a hierarchy of planners to collaboratively solve the scheduling portion of the problem. We envision a high-level planner that produces the concrete workflow graph and delegates to each site’s low-level planner to schedule and connect its assigned components across available interfaces. This approach may prove more efficient for problems with complex resource and/or workflow graphs.